# Mirror Descent Algorithms with Nearly Dimension-Independent Rates for Differentially-Private Stochastic Saddle-Point Problems

### v1

@@ Appendix F Missing details from Section 5
Denote by $G_{j}(S)=\left(\sum_{k\in[i-1]}g^{k}_{j}\right)_{i\in[t]}$ the vector of the cumulative sum of the gradients that our algorithm produces when it has been fed with dataset $S$. Also, define the log-sum-exp function as $\operatorname{LSE}((x_{1},...,x_{k}))=\log(\exp(x_{1})+...+\exp(x_{k}))$. With this new notation
$w^{t}_{j}\propto\exp(\operatorname{LSE}(\tau G_{j}(S))).$
Two well know properties of the $\operatorname{LSE}(x)$ function are that it is convex on its argument and its gradient is the softmax function of $x$. As a consequence of the second property mentioned, it is clear that $\|\nabla\operatorname{LSE}(x)\|_{1}=1$ for every $x$. By convexity of $\operatorname{LSE}$, we have
@@
$\leq\tau\|\nabla\operatorname{LSE}(-\tau G(S))\|_{1}\|G_{j}(S)-G_{j}(S^{\prime})\|_{\infty}$
$=\tau\|G_{j}(S)-G_{j}(S^{\prime})\|_{\infty}.$
Next, note that by $L_{0}$-Lipschitzness of $f$
$\max_{j\in[d_{x}]}\tau\|G_{j}(S)-G_{j}(S^{\prime})\|_{\infty}=\tau\max_{j\in[d_{x}],i\in[t]}\left|\sum_{k\in[i-1]}(g^{k}_{j}-g^{\prime k}_{j})\right|\leq\frac{2L_{0}}{B}.$
Hence, sampling vertices from $\Delta_{x}$ according to the probabilities induced by $w^{t}$ is equivalent to sampling with the exponential mechanism with score function $s(S,j)=\operatorname{LSE}(\tau G_{j}(S))$, whose sensitivity is $\frac{2\tau L_{0}}{B}$. Since we sample $K$ times during the first $q$ iterations, then every $q$ iterations and then $K$ times to sparsify the last iterate, the overall number of samplings will be $TK/q+qK+K$. Since each of them is $\frac{4\tau L_{0}}{B}$-DP, then by Advanced Composition of DP (Theorem A.1) we only need that
$\frac{4\tau L_{0}}{B}\leq\frac{\varepsilon}{2\sqrt{2(TK/q+(q+1)K)\log(1/\delta)}}.$
Proof of Theorem 5.1.
Theorem 5.1.
Let $x\in\Delta_{x}$ be a minimizer of $F_{\mathcal{D}}(\cdot)$. From Lemma 5.1 we have
$F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)\leq\frac{1}{T}\left[\operatorname{Regret}_{T}(x)+\sum_{t\in[T]}\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right],$
where $\operatorname{Regret}_{T}(x)\leq\frac{\log(d_{x})}{\tau}+\frac{\tau}{2}\sum_{t\in[T]}\|g^{t}\|_{\infty}^{2}$ by Lemma A.1. Using that $\|g^{t}\|_{\infty}\leq L_{0}$ and taking expectations, we obtain
@@
Let us bound the term $\mathbb{E}\left[\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right]$. For $t\leq q$, we have $g^{t}=\nabla f(\hat{w}^{t};B^{t})$, so
@@
$\leq 2\|\mathbb{E}_{\hat{w}^{t}}[\nabla f(w^{t};B^{t})-\nabla f(\hat{w}^{t};B^{t})]\|_{\infty}$
$\leq\begin{cases}\frac{8L_{1}}{\sqrt{K}}&\text{without second order smoothnes}\\ \frac{4L_{2}}{K}&\text{with second order smoothness}\end{cases}.$
where in the first equality we used that $\nabla f(w^{t};B^{t})$ is an unbiased estimator of $\nabla F_{\mathcal{D}}(w^{t})$ and for the second inequality we used Corollaries D.2 and D.1. Now we consider the case where $t>q$. First, denote by $t_{0}=q\lfloor t/q\rfloor$. That is, $t_{0}$ was the last time that $\hat{w}$ was updated by the algorithm before the $t$-th iteration. Hence, $g^{t}=\nabla f(\hat{w}^{t_{0}};B^{t})$. Now, we decompose the error into three terms and bound each of them individually:
@@
Clearly $\mathbb{E}[\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(w^{t};B^{t}),x^{t}-x\rangle]=0$, since $\nabla f(w^{t};B^{t})$ is an unbiased estimator of $\nabla F_{\mathcal{D}}(w^{t})$ conditioned on the the randomness up to iteration $t$. For the second term, note that
@@
$\leq 2L_{1}\|w^{t}-w^{t_{0}}\|_{1}\leq 2L_{1}\sum_{k=t_{0}}^{t-1}\|w^{k+1}-w^{k}\|_{1}$
$\leq 2L_{1}\sum_{k=t_{0}}^{t-1}\frac{2}{k+1}\leq\frac{4L_{1}q}{q\lfloor t/q\rfloor+1},$
@@
where the first inequality follows from Hölder, the second from the $L_{1}$-smoothness of $f$, the third from the triangle inequality for the $\ell_{1}$ norm and the fourth from the stability of the average iterates. Now, let us bound the third term: $\langle\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-x\rangle$. We will further decompose this term into
$\langle\tilde{\Delta}^{t_{0}},x^{t}-x\rangle=\underbrace{\langle\tilde{\Delta}^{t_{0}},\tilde{x}^{t}-x\rangle}_{=:A_{1}}+\underbrace{\langle\tilde{\Delta}^{t_{0}},x^{t}-\tilde{x}^{t}\rangle}_{=:A_{2}},$
@@
where $\tilde{\Delta}^{t_{0}}$ denotes $\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})$ and $\tilde{x}^{t}$ is a ficticious iterate obtained by starting at $x^{t_{0}}$ and performing updates $\tilde{x}_{j}^{t+1}\propto\tilde{x}_{j}^{t}\exp\big{(}-\tau\nabla_{j}f(w^{t_{0}};B^{k})\big{)}$. We start by bounding $A_{1}$:
$\mathbb{E}_{\hat{w}^{t_{0}}}[\langle\tilde{\Delta}^{t_{0}},\tilde{x}^{t}-x\rangle]$ $=\langle\mathbb{E}_{\hat{w}^{t_{0}}}[\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})],\tilde{x}^{t}-x\rangle$
$\leq 2\|\mathbb{E}_{\hat{w}^{t_{0}}}[\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})]\|_{\infty}$
@@
$\leq\begin{cases}\frac{8L_{1}}{\sqrt{K}}&\text{without second order smoothnes}\\ \frac{4L_{2}}{K}&\text{with second order smoothness}\end{cases}.$
@@
where the equality holds since $\tilde{x}^{t}$, as opposed to $x^{t}$, does not depend on $\hat{w}^{t_{0}}$, and the last inequality follows from Corollaries D.2 and D.1. By the law of iterated expectation,
$\mathbb{E}[A_{1}]=\mathbb{E}[\langle\tilde{\Delta}^{t_{0}},\tilde{x}^{t}-x\rangle]\leq\begin{cases}\frac{8L_{1}}{\sqrt{K}}&\text{without second order smoothnes}\\ \frac{4L_{2}}{K}&\text{with second order smoothness}\end{cases}.$
We bound now $A_{2}$. Note that $\langle\tilde{\Delta}^{t_{0}},x^{t}-\tilde{x}^{t}\rangle\leq\|\tilde{\Delta}^{t_{0}}\|_{\infty}\|x^{t}-\tilde{x}^{t}\|_{1}$. Note that for $j\in[d_{x}]$
$(x^{t}-\tilde{x}^{t})_{j}=\frac{x^{t_{0}}_{j}\exp\left(-\tau\sum_{k=t_{0}}^{t-1}\nabla_{j}f(\hat{w}^{t_{0}};B^{k})\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(-\tau\sum_{k=t_{0}}^{t-1}\nabla_{i}f(\hat{w}^{t_{0}};B^{k})\right)}-\frac{x^{t_{0}}_{j}\exp\left(-\tau\sum_{k=t_{0}}^{t-1}\nabla_{j}f(w^{t_{0}};B^{k})\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(-\tau\sum_{k=t_{0}}^{t-1}\nabla_{i}f(w^{t_{0}};B^{k})\right)}.$
Denote $\hat{G}_{j}=-\tau\sum_{k=t_{0}}^{t-1}\nabla_{j}f(\hat{w}^{t_{0}};B^{k}),G_{j}=-\tau\sum_{k=t_{0}}^{t-1}\nabla_{j}f(w^{t_{0}};B^{k})$. We rewrite the equality above in terms of $\hat{G}_{j}$ and $G_{j}$ and bound $(x^{t}-\tilde{x}^{t})_{j}$ as follows
@@
$(x^{t}-\tilde{x}^{t})_{j}$ $=\frac{x^{t_{0}}_{j}\exp\left(\hat{G}_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(\hat{G}_{i}\right)}-\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}$
$=\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)\exp\left(\hat{G}_{j}-G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)\exp\left(\hat{G}_{i}-G_{i}\right)}-\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}$
$\leq\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\left[\exp\left(2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\right)-1\right],$
and similarly
$(\tilde{x}^{t}-x^{t})_{j}$ $=\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}-\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)\exp\left(\hat{G}_{j}-G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)\exp\left(\hat{G}_{i}-G_{i}\right)}$
$\leq\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\left[1-\exp\left(-2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\right)\right]$
$\leq\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\left[2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\right].$
Combining both inequalities, we obtain
$|(x^{t}-\tilde{x}^{t})_{j}|\leq\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\underbrace{\max\left\{2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|,\exp\left(2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\right)-1\right\}}_{:=\psi},$
and hence
$\|x^{t}-\tilde{x}^{t}\|_{1}=\sum_{j\in[d_{x}]}|(x^{t}-\tilde{x}^{t})_{j}|\leq\sum_{j\in[d_{x}]}\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\psi=\psi.$
@@
Next, we use that $e^{x}-1\leq 2x$ for $x\leq 1$. Assume for now $2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\leq 1$, so $\psi=4\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|$. Using the definition of $\hat{G}_{j},G_{j}$, it follows that
$\|x^{t}-\tilde{x}^{t}\|_{1}\leq\psi$ $=4\max_{j\in[d_{x}]}\tau\left|\sum_{k=t_{0}}^{t-1}[\nabla_{j}f(\hat{w}^{t_{0}};B^{k})-\nabla_{j}f(w^{t_{0}};B^{k})]\right|$
$\leq 4\tau\max_{j\in[d_{x}]}\sum_{k=t_{0}}^{t-1}\left|\nabla_{j}f(\hat{w}^{t_{0}};B^{k})-\nabla_{j}f(w^{t_{0}};B^{k})\right|$
$\leq 4\tau\sum_{k=t_{0}}^{t-1}\left\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\right\|_{\infty}.$
@@
This has two implications. The first is that $\psi\leq 8\tau qL_{0}$. This further implies that in order to ensure that our assumption $2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\leq 1$ holds, it suffices that $8\tau qL_{0}\leq 2$, or equivalently $\tau\leq 1/(4L_{0}q)$. We impose this condition on $\tau$ later when optimizing the convergence rate. The second is the following bound on $A_{2}$
$A_{2}$ $=\langle\tilde{\Delta}^{t_{0}},x^{t}-\tilde{x}^{t}\rangle$
$=\langle\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-\tilde{x}^{t}\rangle$
@@
$\leq 4\tau\sum_{k=t_{0}}^{t-1}\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}$
$\leq 2\tau\sum_{k=t_{0}}^{t-1}\Big{(}\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}^{2}+\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}^{2}\Big{)}$
$\leq 2\tau q\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}^{2}+2\tau\sum_{k=t_{0}}^{t-1}\Big{(}\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}^{2}\Big{)}$
$\lesssim\begin{cases}q\tau\big{[}\frac{L_{1}^{2}}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}}{\sqrt{K}}\big{]}&\text{without second order smoothnes}\\ q\tau\big{[}\frac{L_{2}^{2}}{K^{2}}+\frac{L_{1}^{2}\log(d_{x})}{K}\big{]}&\text{with second order smoothness}\end{cases},$
where we bounded the second moment of the estimation errors with the results in Corollaries D.3 and D.4. We conclude that for all $t>q$

### v2

@@ 4.1 Privacy Analysis of Algorithm 5
Denote by $G_{j}(S)=\big(\sum_{k\in[i-1]}g^{k}_{j}\big)_{i\in[t]}$ the vector of the cumulative sum of the gradients that our algorithm produces when it has been fed with dataset $S$. Also, define the log-sum-exp function as $\operatorname{LSE}((x_{1},...,x_{k}))=\log(\sum_{i=1}^{k}\exp\{x_{i}\})$. With this new notation, $w^{t}_{j}\propto\exp(\operatorname{LSE}(-\tau G_{j}(S)))$. Two well-known properties of the $\operatorname{LSE}(x)$ function are that it is convex in its argument and its gradient is the softmax function of $x$. As a consequence of the second property mentioned, it is clear that $\|\nabla\operatorname{LSE}(x)\|_{1}=1$ for every $x$. By convexity of $\operatorname{LSE}$, we have
@@
$\leq\tau\|\nabla\operatorname{LSE}(-\tau G(S))\|_{1}\|G_{j}(S)-G_{j}(S^{\prime})\|_{\infty}=\tau\|G_{j}(S)-G_{j}(S^{\prime})\|_{\infty}.$
Furthermore, $\max_{j\in[d_{x}]}\tau\|G_{j}(S)-G_{j}(S^{\prime})\|_{\infty}=\tau\max_{j\in[d_{x}],i\in[t]}\left|\sum_{k\in[i-1]}(g^{k}_{j}-g^{\prime k}_{j})\right|\leq\frac{2\tau L_{0}}{B}$ by $L_{0}$-Lipschitzness of $f$. Hence sampling vertices from $\Delta_{x}$ according to the probabilities induced by $w^{t}$ is equivalent to running the exponential mechanism with score function $s(S,j)=\operatorname{LSE}(-\tau G_{j}(S))$, whose sensitivity is bounded by $\frac{2\tau L_{0}}{B}$. Since we sample $K$ times during the first $q$ iterations, then every $q$ iterations and then $K$ times to sparsify the last iterate, the overall number of samplings will be $TK/q+qK+K$. Since each of them is $\frac{4\tau L_{0}}{B}$-DP, then by Advanced Composition of DP (Corollary 3.6) we only need that $\frac{4\tau L_{0}}{B}\leq\frac{\varepsilon}{2\sqrt{2(TK/q+(q+1)K)\log(1/\delta)}}$.
4.2 Convergence Analysis of Algorithm 5
Proof 4.5 (Proof of Theorem 4.2).
Let $x\in\Delta_{x}$ be a minimizer of $F_{\mathcal{D}}(\cdot)$. From Lemma 4.1, $F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)\leq\frac{1}{T}\big[\sum_{t\in[T]}\langle g_{t},x_{t}-x\rangle+\sum_{t\in[T]}\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\big]$. The standard analysis of mirror descent yields, $\sum_{t\in[T]}\langle g_{t},x_{t}-x\rangle\leq\frac{\log(d_{x})}{\tau}+\frac{\tau}{2}\sum_{t\in[T]}\|g^{t}\|_{\infty}^{2}$ [JNLS:2009]. Using that $\|g^{t}\|_{\infty}\leq L_{0}$ and taking expectations, we obtain
@@ 4.2 Convergence Analysis of Algorithm 5
Let us bound $\mathbb{E}\left[\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right]$. For $t\leq q$, we have $g^{t}\!=\!\nabla f(\hat{w}^{t};B^{t})$, and
@@
$\leq 2\|\mathbb{E}_{\hat{w}^{t}}[\nabla f(w^{t};B^{t})-\nabla f(\hat{w}^{t};B^{t})]\|_{\infty}\leq\begin{cases}\frac{8L_{1}}{\sqrt{K}}&\text{for first-order-smooth $f(\cdot;z)$}\\ \frac{4L_{2}}{K}&\text{for second-order-smooth $f(\cdot;z)$.}\end{cases}$
where in the first equality we used that $\nabla f(w^{t};B^{t})$ is an unbiased estimator of $\nabla F_{\mathcal{D}}(w^{t})$ and for the second inequality we used Corollaries 2.6 and 2.4. Now consider the case $t>q$. First, denote by $t_{0}=q\lfloor t/q\rfloor$. That is, $t_{0}$ was the last time that $\hat{w}$ was updated by the algorithm before the $t$-th iteration. Hence, $g^{t}=\nabla f(\hat{w}^{t_{0}};B^{t})$. Now, we decompose the error into three terms and bound each of them individually:
@@
Clearly $\mathbb{E}[\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(w^{t};B^{t}),x^{t}-x\rangle]=0$, since $\nabla f(w^{t};B^{t})$ is an unbiased estimator of $\nabla F_{\mathcal{D}}(w^{t})$ conditionally on the randomness up to iteration $t$. For the second term, note that
@@
$\leq 2L_{1}\|w^{t}-w^{t_{0}}\|_{1}\leq 2L_{1}\sum_{k=t_{0}}^{t-1}\|w^{k+1}-w^{k}\|_{1}\leq 2L_{1}\sum_{k=t_{0}}^{t-1}\frac{2}{k+1}\leq\frac{4L_{1}q}{q\lfloor t/q\rfloor+1},$
@@
where the first inequality follows from Hölder, the second from the $L_{1}$-smoothness of $f$, the third from the triangle inequality for $\|\cdot\|_{1}$, and the fourth from the stability of the average iterates. Now, we decompose
$\langle\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-x\rangle=A_{1}+A_{2},$
@@
with $A_{1}:=\langle\widetilde{\Delta}^{t_{0}},\tilde{x}^{t}-x\rangle$ and $A_{2}:=\langle\widetilde{\Delta}^{t_{0}},x^{t}-\tilde{x}^{t}\rangle$, where $\widetilde{\Delta}^{t_{0}}$ denotes $\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})$ and $\tilde{x}^{t}$ is a ficticious iterate obtained by starting at $x^{t_{0}}$ and performing updates $\tilde{x}_{j}^{t+1}\propto\tilde{x}_{j}^{t}\exp\big(-\tau\nabla_{j}f(w^{t_{0}};B^{k})\big)$. We start by bounding $A_{1}$:
$\mathbb{E}_{\hat{w}^{t_{0}}}[\langle\widetilde{\Delta}^{t_{0}},\tilde{x}^{t}-x\rangle]=\langle\mathbb{E}_{\hat{w}^{t_{0}}}[\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})],\tilde{x}^{t}-x\rangle$
@@
$\leq 2\|\mathbb{E}_{\hat{w}^{t_{0}}}[\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})]\|_{\infty}\leq\begin{cases}\frac{8L_{1}}{\sqrt{K}}&\text{for first-order-smooth $f(\cdot;z)$}\\ \frac{4L_{2}}{K}&\text{for second-order-smooth $f(\cdot;z)$,}\end{cases}$
@@
where the equality holds since $\tilde{x}^{t}$, as opposed to $x^{t}$, does not depend on $\hat{w}^{t_{0}}$, and the last inequality follows from Corollaries 2.6 and 2.4. We bound now $A_{2}$. Note that $\langle\widetilde{\Delta}^{t_{0}},x^{t}-\tilde{x}^{t}\rangle\leq\|\widetilde{\Delta}^{t_{0}}\|_{\infty}\|x^{t}-\tilde{x}^{t}\|_{1}$. Denote $\hat{G}_{j}=-\tau\sum_{k=t_{0}}^{t-1}\nabla_{j}f(\hat{w}^{t_{0}};B^{k}),G_{j}=-\tau\sum_{k=t_{0}}^{t-1}\nabla_{j}f(w^{t_{0}};B^{k})$. We bound $|(x^{t}-\tilde{x}^{t})_{j}|$ as follows
$(x^{t}-\tilde{x}^{t})_{j}$ $=\frac{x^{t_{0}}_{j}\exp(\hat{G}_{j})}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp(\hat{G}_{i})}-\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}$
$=\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)\exp(\hat{G}_{j}-G_{j})}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)\exp(\hat{G}_{i}-G_{i})}-\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}$
$\leq\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp(G_{i})}\Big[\exp\big(2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\big)-1\Big],$
@@
and similarly $(\tilde{x}^{t}-x^{t})_{j}\leq\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\left[2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\right]$. Combining both inequalities, we obtain $|(x^{t}-\tilde{x}^{t})_{j}|\leq\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\psi$, where
@@
$\psi:=\max\Big\{2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|,\exp\big(2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\big)-1\Big\}$. It follows that $\|x^{t}-\tilde{x}^{t}\|_{1}=\sum_{j\in[d_{x}]}|(x^{t}-\tilde{x}^{t})_{j}|\leq\sum_{j\in[d_{x}]}\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\psi=\psi$. Next, we use that $e^{x}-1\leq 2x$ for $x\leq 1$. Assume for now $2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\leq 1$, so $\psi=4\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|$. Using the definition of $\hat{G}_{j},G_{j}$, it follows that
$\|x^{t}-\tilde{x}^{t}\|_{1}\leq\psi$ $=\textstyle 4\max_{j\in[d_{x}]}\tau\left|\sum_{k=t_{0}}^{t-1}[\nabla_{j}f(\hat{w}^{t_{0}};B^{k})-\nabla_{j}f(w^{t_{0}};B^{k})]\right|$
$\textstyle\leq 4\tau\max_{j\in[d_{x}]}\sum_{k=t_{0}}^{t-1}\left|\nabla_{j}f(\hat{w}^{t_{0}};B^{k})-\nabla_{j}f(w^{t_{0}};B^{k})\right|$
$\textstyle\leq 4\tau\sum_{k=t_{0}}^{t-1}\left\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\right\|_{\infty}.$
@@
This has two implications. The first is that $\psi\leq 8\tau qL_{0}$, by $L_{0}$-Lipschitzness of $f$. This further implies that in order to ensure that $2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\leq 1$ holds, it suffices that $8\tau qL_{0}\leq 2$, or equivalently $\tau\leq 1/(4L_{0}q)$. We impose this condition on $\tau$ later when optimizing the convergence rate. The second is the following bound on $A_{2}$
$A_{2}=\langle\widetilde{\Delta}^{t_{0}},x^{t}-\tilde{x}^{t}\rangle=\langle\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-\tilde{x}^{t}\rangle$
@@
$\textstyle\leq 4\tau\sum_{k=t_{0}}^{t-1}\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}$
$\textstyle\leq 2\tau\sum_{k=t_{0}}^{t-1}\Big(\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}^{2}+\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}^{2}\Big)$
$\textstyle\leq 2\tau q\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}^{2}+2\tau\sum_{k=t_{0}}^{t-1}\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}^{2}$
$\lesssim\begin{cases}q\tau\big[\frac{L_{1}^{2}}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}}{\sqrt{K}}\big]&\text{for first-order-smooth $f(\cdot;z)$}\\ q\tau\big[\frac{L_{2}^{2}}{K^{2}}+\frac{L_{1}^{2}\log(d_{x})}{K}\big]&\text{for second-order-smooth $f(\cdot;z)$,}\end{cases}$
where we bounded the second moment of the estimation errors with the results in Corollaries 2.12 and 2.14. We conclude that for all $t>q$